OpenStudy (anonymous):

Reflect triangle X'Y'Z' over the line y=1 X'=(-11,13), Y'=(-14,17), Z'=(-16,14)

OpenStudy (aum):

The y-coordinate of the first point is 13. y-coordinate of 13 is 12 units ABOVE the line y = 1. When we reflect it over the line y = 1, it will be 12 units BELOW the line y = 1 which makes it 1 - 12 = -11. So reflection of X' (-11,13) over the line y = 1 is (-11, -11). You can do the rest the same way.

OpenStudy (anonymous):

what are the coordinates for X'', Y',' Z''

OpenStudy (aum):

I showed you above X'' = (-11,-11) Do Y'' and Z'' the same way I did X''. The x-coordinate of the point will not change when reflecting over the line y=1. Only the y-coordinate will change. A point (a,b) when reflected over the line y = 1 will change to (a, 2-b) because b is (b-1) above the line y = 1 when reflected, it will be (b-1) below the line. That is, 1 - (b-1) = 1 - b + 1 = 2-b